Two trains, each of length 125 metre, are running in parallel tracks in opposite directions. One train is running in a speed 65 km/hour and they cross each other in 6 seconds. The speed of the other train is
Answer by mananth(16949)

to completely cross each other they have to travel 125+125=250 meters
=0.25 km
let the train speed be x
combined speed = 65+x
6 seconds = 6/3600 = 1/600 hours
t=d/r
0.25/(65+x)=1/600
600*0.25 = 65+x
150=65+x
x=150-65
x=85 mph